4x+ 13= -9 -3(x+9);= -7
Answer by ReadingBoosters(3246)


Distribute the -3
4x + 13 = -9 - 3x - 27 ==> isolate the x variable
4x + 3x = -9 - 27 - 13
7x = -49 ==> divide both sides by 7
x = -7
..............................
Have more questions?
HomeworkHelpers@ReadingBoosters.com.
Delighted to help!
-Reading Boosters
Website: www.MyHomeworkAnswers.com